OpenStudy (anonymous):

I have questions. I need help as soon as possible. I don't want answers, I want help! Please!

OpenStudy (anonymous):

I'll post my questions when someone is willing to help me.

OpenStudy (anonymous):

post first because we need to know what is the question about

OpenStudy (anonymous):

Okay. My fist question is: Solve the system by substitution. -4.5x - 2y = -12.5 3.25x - y = -0.75

OpenStudy (anonymous):

okie i need time you should use simultaneous equation method

OpenStudy (anonymous):

Um. Okay.

OpenStudy (anonymous):

@ganeshie8 could you help when you get the chance? I would really appreciate it so much!

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

do you have answer? to check?

OpenStudy (anonymous):

No I don't even understand how to begin to solve it.

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

Here are my other questions if they come easier to solve for you: Solve the system of inequalities by graphing. \[y \le - 4x - 1 \] \[y > 3x\]

OpenStudy (anonymous):

Sorry on that one it's actually y > 3x - 1

OpenStudy (solomonzelman):

desmos.com google it, it is a graph calculator. You will probably be able to do it, just through desmos....

OpenStudy (anonymous):

okie let me explain....

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

but im not sure about the answer if correct

OpenStudy (anonymous):

@SolomonZelman I have a graphing calculator.

OpenStudy (anonymous):

I'm sorry I'm just getting really frustrated because I'm not getting the help I need and I never do on this site. I try and try, and maybe once someones actually given me math help on this website.

OpenStudy (anonymous):

uhmm @HistoryGeek18 no need sorry but the graph i cant help im weak in that :(

OpenStudy (anonymous):

It's fine. I think I'll just do this by myself and whatever I make on my homework its honestly whatever I'm already failing anyways.

OpenStudy (anonymous):

im sorry cant help strive harder you can do it ^_^